3. (1 pt) The Kap of gold(IIl) hydroxide is 3.0x 10. Show by calculation if a precipitate will form when these two solutions are mixed togcther Solution A: 0.0250 M NaOH (5.00 mL) Solution B: 0.0100 M AuKII) acctate (2.00 ml.) A precipitate will will not Complete the sentence with the correct choice: Cirele oneJust want to verify I'm doing this correctly. Please show all steps, so I can see where I'm going right/wrong. Thank you!
Answer #1

Ksp of Au(OH)3 = 3.0 x 10^-5

concentration of NaOH = 0.0250 x 5 / 5 + 2 = 0.01786 M

concentration of Au3+ = 0.0100 x 2 / 5 + 2 = 0.002857 M

Au(OH)3 --------------> Au3+ +   3 OH-

Q = [Au3+][OH-]^3

   = (0.01786) ( 0.002857)^3

Q = 4.2 x 10^-10

Ksp = 3.0 x 10^-5

here Ksp > Q . so No , precipitate will not form